Systems Staff Meeting Notes 1/4/99

Systems Staff Meeting Notes
January 4, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Suzy Shaw, Ying Tang

Announcements
  1. We have a new employee starting today: Ying Tang is our new Programmer/Analyst.
  2. Telephone Status - voice mail is working. The recent difficulties we've been experiencing seem to be a hardware problem within the system. There is no backup of the current configuration and no printed list of how it is configured. Patrick is currently trying to print out the configuration so he can manually set the system when it is reloaded. The ring service has been changed. The whole system is flaky at the moment. We may have to reset our passwords when the system is reloaded. Bill will add Ying to the system.
Review Minutes of 12/7/98
  1. RIR #1 - Hard drive disks still not received.
  2. Nets #5 - IBM not yet contacted regarding Control Unit leak in MSL.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Three NT work stations have been received. They have gone to Systems personnel: Winston, Suzy & Brian. Winston would like to keep one of the old work stations around.
  2. Bill wants to migrate out Debra & Scott's current machines for the ones replaced by the new NT's.
  3. 12 new Gateway work stations have been received. Need to get a distribution list from Martha.
Networking
  1. Erich Kesse has been having trouble with the Preservation website on the LibPreserve server. He will move it to Smathers NT8.
  2. Closet wiring upgrades - not started yet.
  3. New NT mail server should be up in the next couple of weeks.
  4. File storage messages have been occurring on the current mail server. Storage space is getting full. Winston had a strange incident reported with Lisa Bailey's work station: her storage showed lots of space used, but, in reality, her messages were all small with no attachments. Bill notes that Paul McDonough had a similar problem - he works evenings in the same department - they may be sharing the same work station. Winston will call Paul & investigate.
  5. We've received the go-ahead for new network switches.
Software Projects Status
  1. Mellon - still planning to ship this month as far as we know.
  2. ILL migration to MS SQL: halfway there, full conversion by the end of January.
  3. Licensing database - Brian has been meeting with Marie & Barbara, giving them tutorials on the selector portion of LASSIE.
  4. Newspaper database - first phase is complete. They will use briefcase for remote access. Bill isn't too keen on that idea.
  5. PAULA has been fixed on Bill's machine.
  6. Brian tried to create an application with InterDev on the server - NT 8 wouldn't let him - he needs permission to access.
  7. ATC - need to start preliminary process by next month.
  8. FCLA - migration to NERSP nodes - there were problems, they had to back off. Things are okay now, but the proxy server is not in place yet.
  9. SQL Server 7 - upgrade will be available once we have the upgrade for Back Office suite in place.
DP Coordinator
  1. Students return this week. Expect scheduling to be sporadic until next week.
  2. Debra purged the Compustat Ace Weekly tapes per Compustat's request.
  3. Bill needs to talk to NERDC regarding the problematic Compustat tapes. Will also speak with Peter McKay about renewing the order in an upgraded format.
Calendar
  1. Martin Luther King Day - January 18.
  2. Winston will be out Jan. 25-29.
  3. Debra will be out the afternoon of Jan. 25.
  4. Bill Covey out Jan. 29.
  5. ALA will be Jan. 28 - Feb. 5 in Philadelphia, PA.
  6. Rich will be out to ALA Jan. 29 - Feb. 2.


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu